• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Peter L

    Log entry time 16:15:17 on June26,2004

    Entry number 126986

    keyword=Shift Summary

    So my first shift summary crashed with adaql1...

    We started shift with no beam due to a septum valve malfunction. Mark was brought in and he tried to move the valve through software and this failed resulting in the valve opening 100%. This tripped both Q1's and Mark had to manually control the valve. He came out thinking the ordeal was over, and soon after we watched the coolant levels drop once again. Also the JT valve percent readout on the septum screen read negative. Thus Mark needed to return with a new CPU unit. He installed this and Darcy took the opportunity to take a look at the left scanner (still giving us position readout problems) The septums needed to be ramped down for the CPU replacement so we were down for quite some time. Mark solved the problem and Darcy got out of the hall and we started doing BCM calibration runs... they may not be of any use because the left septum was ramping and the LH2 target was in.

    While the hall was open we took scanner runs hoping to understand the left scanner position discrepencies.

    2427 pedestal
    2428,2429,2430 scanner testing
    *these may not be useful*
    2431 BCM 1uA
    2432 BCM 5uA
    2433 BCM 10uA
    2434 BCM 15uA

    2435 Production Right Septum ramping in the Left
    2436 junk
    2437 Production, adaql1 crashed
    Bryan rebooted adaql1 (first time it checked disks in 245 days!)
    WE took advantage of the down time to move the target to carbon.
    BCM runs listed in Kent's note.

    Note: during target change it stopped. we reset the position but to no avail. so we restarted IOC, I had to manually heat the target while doing so, and it moved after the restart.

    Production now finally and we are going home.